# Combining Multi-Objective Bayesian Optimization with Reinforcement Learning for TinyML

### v2

@@ 5.3. Comparison with MCUNet and Zero-Shot NAS
We present a comparison between our results from Fig. 1, five DNN candidates from the MCUNet model list (Lin et al., 2020, 2021), which are optimized for ImageNet and which we retrained for CIFAR10, and a version of MCUNet that we optimized specifically for CIFAR10 using ProxylessNAS (Cai et al., 2018), see Table 2. The original input resolution of the CIFAR10 dataset is 32x32 which is used by our models. For the MCUNet models taken from the model list we report different input resolutions as they were predetermined. For ProxylessNAS, 48x48 was the smallest possible resolution that did not crash the framework. A comparison with the other time series datasets we present in our evaluation is not easily possible, since both MCUNet and ProxylessNAS do not support input with only one spatial dimension. Since our tool reports FLOPs instead of MACs, we assume that one MAC operation (multiply-accumulate) equals two FLOPs, i.e., a multiplication followed by an addition. Since MCUNet uses a different mapping tool than we do, we also report the SRAM and Flash consumption of the mapping achieved by our tool in parentheses.
@@
We observed that our NAS approach was able to find network candidates that achieve a slightly higher accuracy than all tested MCUNet variants, while offering an overall better memory footprint, especially in terms of ROM consumption, at the cost of being more computationally complex. One reason for the higher computational complexity may be that MCUNet uses computationally less intensive depthwise separable convolutions (Chollet, 2017), while our optimized architecture, i.e. ResNet18, uses regular convolutions. In addition, we noticed that the dynamic first-fit memory allocator of the DNN mapping tool we used often reported significantly lower RAM consumption for smaller input resolutions than MCUNet’s patch-based strategy, while the opposite was true for larger input sizes.
@@
5.4. Intuition
(a) ParEGO
(b) ARS
Figure 3. Topography of Bayesian surrogates (top row) and EI (bottom row) after 40 samples for both ParEGO and ARS (ours) given two sets of priors marked with red triangles on a synthetic minimization problem (Passino, 2005) (ground truth see Appendix C, global minimum at $\theta_{0}=15$, $\theta_{1}=5$). For ARS, the rollouts of the trained policies of the 40th sample are shown as lines with their starting points marked by red crosses.
To better understand why our RL-based solver can improve on other Bayesian solvers like ParEgo, we analyzed both using a synthetic optimization problem. Suppose we minimize a two-dimensional single objective optimization problem (Passino, 2005) using both ParEgo and our approach, see Fig. 3. We chose this problem because it presents an interesting challenge: it has a landscape with large areas with little to no gradient combined with some steep minima and maxima. We provide the topography of the objective function in Appendix C and show the respective Bayesian surrogates of ParEgo and ARS-MOBOpt as well as the topographies of their expected improvement (EI) after 40 trials in Fig. 3(a) and 3(b). For our approach (ARS-MOBOpt), we also show the rollouts performed during the validation of the local policies for five selected start points (red crosses) after the last (40th) sample.
We manually selected two sets of priors (marked as red triangles) to provide both algorithms with interesting starting situations: First, a situation where all samples are close together in a region with almost no gradient, see top left in Fig. 3(a) and 3(b). This initialization is interesting because it provides almost no information about the environment, requiring extensive exploration by the algorithms before they can exploit found minima. Second, a more general starting situation with a broader distribution of prior samples, see top right in Fig. 3(a) and 3(b). This initialization focuses more on evaluating how well the algorithms can apply the exploration-exploitation tradeoff and how fast they can find minima. For both priors, we took care not to include samples close to the global minimum.
Looking at the results proposed by ParEgo (marked as black dots) after 40 samples in Fig. 3(a), it is clear that for both sets of priors the algorithm seems to have a clear focus on exploration, leading to a wide sampling of the search space. However, looking at the two examples, this can lead to situations where even though the global optimum was either not found (left column in Fig. 3(a)) or it was not fully exploited (right column in Fig. 3(a)). Furthermore, we argue that this focus on exploration is mainly a result of the definition of the EI as it does not consider any ”negative improvement”. Looking at the topography of the EI for both examples in the last row of Fig. 3(a), it is clear that this can result in large parts of the acquisition function having no gradient. We argue that this is not optimal when relying on gradient-based solvers, as the ParEgo implementation typically does.
In comparison, if we look at the result of our proposed ARS-based approach in Fig. 3(b), it can be seen that our competing local policies, visualized as lines with red crosses as their starting point, have become experts in solving the environment around their starting point and can therefore provide sound solutions even in regions without significant gradients. Therefore, compared to ParEgo, we observed less undirected exploration and a clearer focus on exploiting environmental knowledge, compared to Fig. 3(a). However, despite this more pronounced focus on exploitation, due to the competing ensemble of agents, our solver was still able to consistently escape local minima, e.g. see Fig. 3(b) (right col.).
@@ 6. Conclusion
We presented a NAS methodology for efficient multi-objective DNN optimization for microcontrollers based on Bayesian optimization and Augmented Random Search (ARS). Since the objectives accuracy, ROM consumption, RAM consumption and FLOPs were expensive to evaluate and we were faced with a limited search budged, we focused on performing time-efficient optimization of both DNN hyper- and compression-parameters to enable an optimal deployment on microcontroller platforms. We provided results for two different problems considering two datasets and DNN architectures and showed empirically that our algorithm was able to yield a better feasible Pareto front compared to well-known Bayesian optimization strategies like ParEGO, TurBO or MorBO.
Acknowledgements.
This work was supported by the Bavarian Ministry for Economic Affairs, Infrastructure, Transport and Technology through the Center for Analytics-Data-Applications (ADA-Center) within the framework of “BAYERN DIGITAL II”.
Appendix A Training Details & Optimization Variables
We present a multi-objective Bayesian optimization approach for DNN hyperparameter optimization that uses deep compression, i.e., pruning and quantization, to find feasible tradeoffs between DNN accuracy and resource consumption. Furthermore, we present a novel solver based on Augmented Random Search. An illustration of our process can be found in Fig. 4, with its implementation based on PyTorch, the Optuna (Akiba et al., 2019) framework and BoTorch (Balandat et al., 2020). The implementation consists of two building blocks: (1) the multi-objective optimization loop marked in blue and (2) a set of objective function evaluators marked in orange, for which we use the implementation discussed in (Deutel et al., 2023). The optimization loop iteratively proposes new sets of hyperparameter configurations using ARS as its solver until a termination condition is met.
Figure 4. Overview of our proposed RL-based multi-objective Bayesian DNN hyperparameter optimization approach for microcontroler targets.
Parameter Search Space Interval
epochs uniform $[100,500]$
batch size uniform $[20,200]$
learning rate (lr) log. uniform $[1e^{-5},1e^{-2}]$
momentum log. uniform $[0.7,0.99]$
lr schedule uniform $[0.4,0.9]$
lr gamma uniform $[0.4,0.9]$
weight decay log. uniform $[0.6,0.99]$
pruning start uniform $[0.0,0.6]$
pruning end uniform $[0.8,0.95]$
pruning steps uniform $[1,20]$
pruning sparsity uniform $[0.1,0.99]$
We provide a complete listing of the hyperparameters that are optimized and their search space intervals in Table 3. The first seven parameters are hyperparameters related to general DNN training, while the last five are related to compression. Note that each trainable layer (convolutions and linear layers) has its own optimizable pruning sparsity, and that we use iterative pruning with the pruning start and pruning end parameters defined relative to the number of training epochs. The search space intervals are chosen either based on expert knowledge, e.g., learning rate, momentum or weight decay, or based on time and hardware constraints of the optimiztion, e.g., training epochs and batch size.
Appendix B Complete Listing for UEA & ICR Time Series Classification Benchmark
In Table 4 we provide all results we obtained for the multivariate datasets of the UEA & UCR time series classification benchmark (Bagnall et al., 2017). We provide results for our ARS strategy (columns 3–7) and ParEGO (columns 8–12). For both strategies, we show the achieved hypervolume after 150 trials as well as the maximum accuracy. We also show the minimum memory and FLOPs required to achieve at least 70% accuracy. If no trained networks were found that met these requirements (marked in gray), we show the memory and FLOPs consumption of the network that achieved the highest accuracy. We provide single objective reference results for InceptionTime (Ismail Fawaz et al., 2020) as taken from (Bagnall et al., 2017).
@@ Appendix B Complete Listing for UEA & ICR Time Series Classification Benchmark
Appendix C Synthetic Minimization Problem

### v3

@@ 5.4.2. Comparison with MCUNet and Zero-Shot NAS
We present a comparison between our results from Fig. 2, five DNN candidates from the MCUNet model list (Lin et al., 2020, 2021), which are optimized for ImageNet and which we retrained for CIFAR10, and a version of MCUNet that we optimized specifically for CIFAR10 using ProxylessNAS (Cai et al., 2018), see Table 3. The original input resolution of the CIFAR10 dataset is 32x32 which is used by our models. For the MCUNet models taken from the model list we report different input resolutions as they were predetermined. For ProxylessNAS, 48x48 was the smallest possible resolution that did not crash the framework. A comparison with the other time series datasets we present in our evaluation is not easily possible, since both MCUNet and ProxylessNAS do not support input with only one spatial dimension. Since our tool reports FLOPs instead of MACs, we assume that one MAC operation (multiply-accumulate) equals two FLOPs, i.e., a multiplication followed by an addition. Since MCUNet uses a different mapping tool than we do, we also report the SRAM and Flash requirements of the mapping achieved by our tool in parentheses.
@@
We observed that our NAS approach was able to find network candidates with slightly higher accuracy than all MCUNet variants tested, while offering an overall better memory footprint, especially in terms of ROM, at the cost of higher computational complexity. One reason for the higher computational complexity may be that MCUNet uses less computationally intensive depthwise-separable convolutions (Chollet, 2017), while our optimized architecture, i.e. ResNet18, uses regular convolutions. In addition, we noticed that the dynamic first-fit memory allocator of the DNN mapping tool we used often reported significantly lower RAM requirements for smaller input resolutions than MCUNet’s patch-based strategy, while the opposite was true for larger input sizes.
@@
5.4.3. Comparison with AMC and Automated Pruning
Figure 6. Optimization of MobileNet and CIFAR10 using ARS-MOBOpt (ours) and AMC. We ran our optimization for 50 samples while AMC by default only delivers a single tradeoff.
[…]
The feasible Pareto sets determined by ARS-MOBOpt for MobileNet and CIFAR10 compared to AMC. AMC produced only a single tradeoff that, while generally comparable in performance, is dominated by several of the feasible tradeoffs proposed by our approach in regards of accuracy, ROM, RAM, and FLOPs.
Another approach to find an optimal pruning schedule for a given DNN architecture, is automated pruning (He et al., 2018; Alwani et al., 2022; Hirsch and Katz, 2022). Most automated pruning algorithms are based on sensitivity analysis (Han et al., 2016). The idea of this method is that, given a well-trained baseline model, the sensitivity of each of its layers, i.e., the percentage of a layer’s neurons that can be removed by pruning before the overall accuracy of the baseline DNN starts to degrade, can be easily queried iteratively layer by layer. Typically, during sensitivity analysis the baseline model is not retrained after pruning, making the method very time efficient. Once a sensitivity has been queried for each layer, a pruning schedule, i.e., a set of pruning rates for each layer, can be derived. This schedule can then be executed in a second training run, resulting in a well-trained and pruned DNN.
In the following, we compare our ARS-MOBOpt approach with AMC (He et al., 2018), an automated pruning approach that uses a DDPG RL agent (Lillicrap et al., 2015) to iteratively query pruning rates for each layer given a well-trained baseline network. We search for pruned versions of MobileNetV1 on CIFAR10 using both our approach in blue and AMC in orange, see Fig. 6. To make the results of our apporach and AMC directly comparable regarding memory and FLOPs, we use the same quantization and deployment method we described in Section 5.1 for both the pruned model generated by AMC and ours. Since AMC proposes only a single pruning schedule, i.e., a single tradeoff, instead of a Pareto front of tradeoffs, only a single point is shown for AMC in Fig. 6.
Considering only accuracy, we observed that AMC found a DNN pruning configuration that produced a DNN model comparable to the best model found by our approach. However, when also considering the other objectives, it can be seen that especially for RAM and FLOPs, our approach found significantly better tradeoffs than AMC. We believe that this is the result of two aspects: First, all approaches based on sensitivity analysis do not consider retraining after pruning. However, it is well known that a significant amount of accuracy initially lost by pruning can be recovered later by retraining (Han et al., 2016; Li et al., 2017; Deutel et al., 2023). Since sensitivity analysis does not consider retraining or pruning multiple layers together, it tends to produce conservative pruning schedules that overestimate the effect pruning has on the DNN. Second, AMC does not explicitly consider memory, only FLOPs in the reward of its DDPG agent, and also does not allow the explicit modeling of constraints. However, while this would explain why our framework found better tradeoffs between accuracy and memory, since, unlike AMC, it explicitly considers memory during optimization, it still does not fully explain why AMC did not find models with a similar number of FLOPs compared to the models proposed by our approach.
@@ 6. Conclusion
We presented a NAS methodology for efficient multi-objective DNN optimization for microcontrollers based on Bayesian optimization, Augmented Random Search (ARS), and pruning. Since the objectives accuracy, ROM, RAM and FLOPs were expensive to evaluate and we were faced with a limited search budged, we focused on performing time-efficient optimization of both DNN hyper- and compression-parameters to enable an optimal deployment on microcontroller platforms. We provided results for two different problems considering two datasets and DNN architectures and showed empirically that our algorithm was able to yield a better feasible Pareto front compared to well-known Bayesian optimization strategies like ParEGO, TurBO or MorBO.
Appendix A Complete Listing for UEA & ICR Time Series Classification Benchmark
In Table 4 we provide all results we obtained for the multivariate datasets of the UEA & UCR time series classification benchmark (Bagnall et al., 2017). We provide results for our ARS strategy (columns 3–7) and ParEGO (columns 8–12). For both strategies, we show the achieved hypervolume after 150 trials as well as the maximum accuracy. We also show the minimum memory and FLOPs required to achieve at least 70% accuracy. If no trained networks were found that met these requirements (marked in gray), we show the memory and FLOPs requirements of the network that achieved the highest accuracy. We provide single objective reference results for InceptionTime (Ismail Fawaz et al., 2020) as taken from (Bagnall et al., 2017).
@@ Appendix A Complete Listing for UEA & ICR Time Series Classification Benchmark
Appendix B Synthetic Minimization Problem
@@
[…]
Topography of the synthetic minimization problem.
Appendix C PPO Hyperparameters
Table 5 shows the PPO hyperparameters found that maximize the Hypervolume improvement of on of the multi-objective optimization problems discussed in Sec. 5.4.1, ResNet trained on CIFAR10, compared to the default hyperparameters [^2] https://stable-baselines3.readthedocs.io/en/master/modules/ppo.html of the algorithm.
Table 5. PPO hyperparameters found that maximize the Hypervolume improvement of ResNet trained on CIFAR10 (Optimized) compared to the default parameters (Default).
Parameter Search Space Interval Default Optimized
batch_size categorical $(32,64,128)$ 64 64
clip_range categorical $(0.1,0.2,0.3,0.4)$ 0.2 0.4
ent_coef uniform $[1e^{-8},0.1]$ 0.0 0.016
gae_lambda categorical $(0.92,0.95,0.98)$ 0.95 0.98
gamma categorical $(0.9,0.95,0.98,0.99,0.995)$ 0.99 0.98
learning_rate uniform $[1e^{-5},1]$ $3e^{-4}$ $5e^{-4}$
max_grad_norm categorical $(0.3,0.5,0.6,0.7,0.8,0.9,1,2,5)$ 0.5 1.0
n_epochs categorical $(5,10,20)$ 10 20
n_steps categorical $(16,32,64,128)$ 2048 32
vf_coef uniform $[0,1]$ 0.5 0.24